A body of mass m is suspended by two strings making angles alpha and beeta with the horizontal. Find tensions in the strings.

Grade:11

7 Answers

sagar kumar
6 years ago
T1cos alpha=T2cos beeta--eqn1T1sin alpha+T2sin beeta =mg--eqn2Firstly from eqn 1 obtain T2In eqn2 put the value of T2 and obtain T1And then again put the value of T1 in T2Then finally you will get answer
Shailendra Kumar Sharma
6 years ago
So T1 makes alpha and T2makes beta then vertical components balance will giveT1sin(alpha) + T2sin(beta) = mgHorizontal components T1cos(alpha) = T2cos(beta)T2= T1cos(alpha). / cos (beta)T1sin(alpha)+T2 tan( beta)*cos(alpha)=mgT1=mgsec(alpha)/(tan(alpha) + tan(beta)
Kuldip Pawar
6 years ago
It`s solve by Levi`s Theorem:-->T1/sin(90+β)=T2/sin(90+α)=TE/sin(180-sinβ+sinα).Then---->T1/cosβ=T2/cosα=T3/cos(β+α).
Gitanjali Rout
5 years ago
T1cos alpha=T2cos beeta--eqn1T1sin alpha+T2sin beeta =mg--eqn2Firstly from eqn 1 obtain T2In eqn2 put the value of T2 and obtain T1And then again put the value of T1 in T2Then finally you will get answer//,.k
Gitanjali Rout
5 years ago
o T1 makes alpha and T2makes beta then vertical components balance will giveT1sin(alpha) + T2sin(beta) = mgHorizontal components T1cos(alpha) = T2cos(beta)T2= T1cos(alpha). / cos (beta)T1sin(alpha)+T2 tan( beta)*cos(alpha)=mgT1=mgsec(alpha)/(tan(alpha) + tan(beta)
Gitanjali Rout
5 years ago
o T1 makes alpha and T2makes beta then vertical components balance will giveT1sin(alpha) + T2sin(beta) = mgHorizontal components T1cos(alpha) = T2cos(beta)T2= T1cos(alpha). / cos (beta)T1sin(alpha)+T2 tan( beta)*cos(alpha)=mgT1=mgsec(alpha)/(tan(alpha) + tan(beta)hlkj ….........
Forum Team
5 years ago
 
 
Gitanjali, please stop copy pasting answers posted by other users. You will not be rewarded for these points and your account will be blocked due to this activity as it is against forum guidelines. Please check the guidelines here – Forum point Policy – https://www.askiitians.com/forum-point-policy/
 
Thanks
Forum Team